OpenStudy (anonymous):

Please help me :D Problem: Two cars A and B are 200km apart and they are traveling toward each other at the speeds of 28km/h and 22 km/h respectively. When and where will they meet? 2. If the car A and B in problem 1 start from the same point, going toward the same direction, at what time will they be 18km apart? 3. If B starts at 1:00pm and gives chase starting at 1:15 from the same point as B, when and where will A overtake B? Thank you so much for answering!!

OpenStudy (anonymous):

1) relative speed =28+22=50 they will meet after 200/50=4 hr at a distance of 28*4=112 km from the starting poing or (200-112=88 from the other sid)

OpenStudy (anonymous):

2) relative speed now =28-22=6 they will be 18 km apart after 18/6 = 3 hr

OpenStudy (anonymous):

where did you get 28?

OpenStudy (anonymous):

s it 3) in 15min b has already gone 22/4 =11/2 km their relative speed =28-22=6km/hr time required to overtake is (11/2)/6 =11/12 hr =11/12 * 60 =55 mins(after 1:15) i.e at 2:10 and distance from the starting point is 7/6*22 km =77/3 km the speed of A is 28 km/hr as mentioned

OpenStudy (anonymous):

what formula did you use? I really don't know how you did that??????????????

OpenStudy (anonymous):

disatnce = speed x time

OpenStudy (anonymous):

if two bodies move in same direction their relative speed =difference of their individual speed if they move in same direction then relative speed is the sum of their individual speeds

OpenStudy (anonymous):

ok... proceed

OpenStudy (lgbasallote):

@alynamarie do you agree that distance = speed x time? it is derived from speed = distance/time if you cross multiply you get distance = speed x time

OpenStudy (anonymous):

yes igbasallote

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

LGBASALLOTE* and do you agree that the distance traveled by car A + distance traveled by car B = total distance between them?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

uhmm wait..i read the question wrong..

OpenStudy (anonymous):

it's alright

OpenStudy (anonymous):

Careful....

OpenStudy (lgbasallote):

i thought it was going away...it was towards

OpenStudy (lgbasallote):

so let's start from the beginning heh

OpenStudy (anonymous):

Let us start from the start.. He he.

OpenStudy (anonymous):

sure :)

OpenStudy (lgbasallote):

uhh feeling repetitive @waterineyes ? lol

OpenStudy (anonymous):

Ha ha ha..

mathslover (mathslover):

Dont want to break the flow of the answers but just a notification for @alynamarie ; Please post right questions in right group ... This is a physics question , so please post it in physics section For further details : http://openstudy.com/code-of-conduct

OpenStudy (lgbasallote):

actually this can be classified as math as well

OpenStudy (lgbasallote):

since it just uses the basic speed = distance/time

OpenStudy (anonymous):

Actually @lgbasallote is also right..

OpenStudy (lgbasallote):

-_-

mathslover (mathslover):

Physics can be fully classified as math.. There is calculation is every where in physics that is basically maths BUT THEN WHY WE HAVE PHYSICS SECTION RIGHT UP HERE .. Please obey the code of conduct instead of having arguments against it .. Thanks!

OpenStudy (lgbasallote):

@lgbasallote isnt *also* right... @lgbasallote *is* right -_- @waterineyes

mathslover (mathslover):

Don't spam! Common forms of spam include: Posting a question that is off topic for a particular group (i.e. posting a physics question in the Mathematics group). -Code of Conduct

OpenStudy (lgbasallote):

@mathslover this is mathematics curriculum NOT physics

mathslover (mathslover):

Ok so! it is maths ... I didn't know about that @lgbasallote So the physics section and code of conduct are just for enjoyment right ?

OpenStudy (lgbasallote):

alright...you're the ambassador..you know better

OpenStudy (anonymous):

dear all i think this is a common topic for maths as well as physics...

mathslover (mathslover):

@alynamarie by the way did u get the answer ?

OpenStudy (anonymous):

@mathslover you can post this question in physics as well as Maths Section.. It is not against the code of conduct..

OpenStudy (anonymous):

yup, this is math

mathslover (mathslover):

I will suppose to continue this discussion with a mod instead .@waterineyes and @all.. BY the way u can continue for the solutions

OpenStudy (anonymous):

@mathslover nope, haven't got the answer yet

mathslover (mathslover):

Should i help u now ?

OpenStudy (anonymous):

@mathslover @lgbasallote is helping someone and you just came and you stopped someone to learn.. Couldn't you do this discussion after the question has been solved??? What is the problem in that..

OpenStudy (anonymous):

please help me... anyone

OpenStudy (anonymous):

even if some body asks other subjects questions idont think it is wrong the important thing has to be the question asked has to be answered properly...

OpenStudy (anonymous):

The question is of maths and physics both.. Anyone can post it where he wants..

OpenStudy (anonymous):

yeah... could someone please help me answer and understand this?? Please :D

mathslover (mathslover):

Two cars A and B are 200km apart and they are traveling toward each other at the speeds of 28km/h and 22 km/h respectively. When and where will they meet? Given distance = 200 km. u = 22 km/h v = 28 km/h Let that they will meet at point h.. Hence one part will be : 200 - h and other will be h

mathslover (mathslover):

I am trying this @alynamarie please wait

OpenStudy (anonymous):

ok @mathslover

OpenStudy (anonymous):

maybe it's better to solve this in terms of mathematics like their notations. u and v are somewhat physics looking

mathslover (mathslover):

the process that is shown by @matricked is very right

OpenStudy (anonymous):

i know. but i just don't know the formula used

mathslover (mathslover):

http://en.wikipedia.org/wiki/Relative_velocity

mathslover (mathslover):

\[\vec{v}_{BA}=-\vec{v}_{AB}\]

OpenStudy (anonymous):

ok. that's it 'm closing this yet reposting it in the physics group..... anyone willing to start over :) pleaseee

mathslover (mathslover):

k thanks .. that is what i was wanting! thanks for ur support

OpenStudy (anonymous):

but no one is answering in the physics section

OpenStudy (anonymous):

see what i mean..... no one is answering my quetion: posted in the physics section :(

mathslover (mathslover):

tag people dude @--- like @mathslover

OpenStudy (asnaseer):

@mathslover I agree with your general sentiment that questions should indeed be posted in the right group so that others in that groups can also benefit from it. However, in this particular instance the question is really in a /grey/ area between maths and physics. So I believe it was fine for this to be posted here.

mathslover (mathslover):

k thanks sir

OpenStudy (asnaseer):

yw :)